OpenStudy (anonymous):

Line AB contains points A (3, -2) and B (1, 8). The slope of line AB is

OpenStudy (anonymous):

Select one: a. -5 b. -1/5 c. 1/5 d. 5

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@purplepiggy171

OpenStudy (cggurumanjunath):

(y2-y1)/(x2-x1)=slope (x1,y1)=(3,-2) (x2,y2)=(1,8)

OpenStudy (cggurumanjunath):

@prowrestler

OpenStudy (anonymous):

slope = -5

OpenStudy (anonymous):

so ur saying the answer is a

OpenStudy (anonymous):

Yeah..

OpenStudy (anonymous):

8-(-2)/1-3 = -5 right?

OpenStudy (cggurumanjunath):

yes slope=-5

OpenStudy (triciaal):

slope is change in y divided by change in x yes -5

OpenStudy (anonymous):

thanks